# Hyperdimensional Computing for Node Classification and Link Prediction

### v2

@@ Abstract
We introduce a novel method for transductive learning on graphs using hyperdimensional representations. The proposed approach encodes data samples using random projections into a very high-dimensional space (hyperdimensional or HD space for short). It obviates the need for expensive iterative training of the sort required by deep learning methods. Specifically, we propose a Hyperdimensional Graph Learning (HDGL) algorithm. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
We report results of experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost. Furthermore, HDGL is well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes. Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods, although it falls short of state-of-the-art Graph Neural Network (GNN) methods that rely on computationally expensive iterative training. We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.
1 Introduction
@@
Node Classification and Link Prediction in Networks. Social networks, citation networks, etc., are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities. For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations. Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction [30, 29]. Such problems are typically solved using graph embedding techniques [63, 6, 60, 55].
1.1 Related Work
@@
Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita [46], Frasconi et al. [13], Gori et al. [17], Scarselli et al. [44], which learn low-dimensional embeddings of nodes in a graph by aggregating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning [26, 53, 18], because of their superior performance across a broad spectrum of applications [58, 62, 63]. Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact [10, 47, 9, 50]. Hence, there is an urgent need for computationally more efficient, and consequently, environmentally more sustainable alternatives to deep learning models in general, with a specific emphasis on graph neural networks.
@@
Hyperdimensional computing [22, 23, 37, 49] offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector [22, 23]. All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products [22, 37, 49]. Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware [27, 28]. Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning [37, 28, 15, 34].
@@
Hyperdimensional encodings of graphs have been investigated by Poduval et al. [43] and Nunes et al. [38]. These works focus on performing graph operations, e.g., similarity based retrieval, graph classification, etc. In contrast, our focus is on methods that learn node-level representations for node classification and link prediction.
Hyperdimensional node embeddings have been explored by Kang et al. [24] and [31]. Specifically, Kang et al. [24] proposed RelHD, a novel processing-in-memory (PIM) hardware architecture [35] based on FeFET technology [4]; and Li et al. [31] introduced HyperNode for learning HD embeddings of nodes. However, both RelHD and HyperNode suffer from significant shortcomings associated with how they construct node representations from node features (see section 3.5 for details).
1.2 Overview and Contributions
We introduce a novel method for transductive learning on graphs using hyperdimensional representations. The proposed approach encodes data samples using random projections into a very high- dimensional space (hyperdimensional or HD space for short) [22, 23, 37, 49]. It obviates the need for expensive iterative training of the sort required by deep learning methods. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
Contributions. The key contributions of the paper are as follows:
  • We introduce HDGL, a method for learning HD node embeddings that match the expressive power of GNN embeddings. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
  • We show how the HD node embeddings learned by HDGL can be used to perform both semi-supervised node labeling and link prediction tasks. Thus, unlike deep learning methods usually require separate models for each of these tasks, HDGL produces a single model that can support both tasks.
  • We present results of extensive experimental comparisons of HDGL with the state-of-the-art GNN methods on several standard benchmark data sets which show that HDGL achieves performance that is competitive with that of GNN, while avoiding the need for computationally expensive training.
  • We present results experiments which show that HDGL offers a better alternative to graph neural networks for settings that call for class-incremental learning, where the model is required to effectively discriminate between a growing number of classes that are presented incrementally over time.
Organization. The rest of the paper is organized as follows: Section 3 reviews the HD computing framework, its basic operations, and their properties, and reviews the expressive power of GNN. Section 4 introduces HDGL, the proposed HD computing based algorithm for learning node embeddings. Section 5 describes results of extensive experiments that compare HDGL with the state-of-the-art GNN-based node embedding methods. Section 6 concludes with a brief summary of the paper and an outline of some directions for further research.
2 Problem Statement
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of undirected nodes, $\mathcal{E}$ represents the set of edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$. Let $N=|\mathcal{V}|$ denote the number of nodes in the graph. Supposed each node $v\in\mathcal{V}$ is described by a tuple of $d$ features $\mathbf{x}_{v}\in\mathbb{R}^{d}$. The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$.
For the node label prediction problem in the transductive setting, let $y_{w}\in{1,\cdots,L}$ represent the label associated with each node $w\in\mathcal{W}$, where $\mathcal{W}\subset\mathcal{V}$. This implies that not all nodes in the graph are labeled. The goal is to learn a latent representation $\phi(v)$ for each node $v$, utilizing the structure of $\mathcal{G}$, the feature vector $\mathbf{x}_{v}$, and the labels $y_{w}$, to accurately predict the labels for a set of nodes that are unlabeled in the graph $\mathcal{G}$.In the transductive link prediction setting, the goal is to infer missing edges within the graph by leveraging the latent representations $\phi(v)$ of the nodes.
3 Background and Preliminaries
3.1 Properties of HD Computing
@@
Hyper-dimensional computing, originally introduced by Kanerva [22] is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large [22, 23]. All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products [22, 37]. The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations [27, 28]. In this paper, we assume that =$\mathcal{H}=\{0,1\}^{n}$. We proceed to briefly summarize some of the key properties of HD representation and operations that can be used to compute with it.
@@
Near Orthogonality of Random vectors in HD Space. HD encodings of objects are produced by sampling each dimension independently and uniformly from a distribution. Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$. We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ [22]. This is also known as a blessing of HD spaces.
@@
High Memory Capacity of HD Representation. As noted by Neubert et al. [37], in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$. For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
@@
Noise or Error Tolerance. Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits [1, 37, 49]. Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms [37].
3.2 HD Computing Operations
@@
Bundling, denoted by $\oplus$, is a bitwise majority operation represented by :$\oplus:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. Bundling two randomly sampled HD vectors results in a new vector which is similar to the input vectors. Bundling $\oplus$ is associative and commutative i.e $\mathbf{a}\oplus(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\oplus\mathbf{b})\oplus\mathbf{c}=(\mathbf{c}\oplus\mathbf{a})\oplus\mathbf{b}$. Note that in case of Bundling an even number of HD vectors, there is a possibility of ties since Bundling is a bitwise majority operation. To break ties, a tie-breaking policy, e.g., random choice of 0 or 1 on dimensions where ties occur. Bundling can be thought of as the HD space analog of the mean operation euclidean spaces. Bundling, because it commutes, can be used to represent sets as well as multisets of objects encoded using HD vectors.
@@
Rotation, also called permutation, denoted by $\Pi$, is a bitwise left shift operation $\Pi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$. Rotation was introduced by Gayler [14] in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$. In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
@@
3.3 GNN-based Node Embdeddings

### v3

# Hyperdimensional Representation Learning for Node Classification and Link Prediction

@@ Abstract.
We introduce Hyperdimensional Graph Learner (HDGL), a novel method for node classification and link prediction in graphs. HDGL maps node features into a very high-dimensional space (hyperdimensional or HD space for short) using the injectivity property of node representations in a family of Graph Neural Networks (GNNs) and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node yielding latent node representations that can support both node classification and link prediction tasks. HDGL, unlike GNNs that rely on computationally expensive iterative optimization and hyperparameter tuning, requires only a single pass through the data set. We report results of experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL achieves accuracy that is competitive with that of the state-of-the-art GNN methods at substantially reduced computational cost; and on the link prediction task, HDGL matches the performance of DeepWalk and related methods, although it falls short of computationally demanding state-of-the-art GNNs.
Hyperdimensional Computing, Graph Neural Networks, Transductive Learning, Representation Learning
journalyear: 2025copyright: acmlicensedconference: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanybooktitle: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining (WSDM ’25), March 10–14, 2025, Hannover, Germanydoi: 10.1145/3701551.3703492isbn: 979-8-4007-1329-3/25/03ccs: Computing methodologies Learning latent representationsccs: Computing methodologies Semi-supervised learning settings
1. Introduction
@@
Social networks, citation networks, molecular, e.g., protein-protein interaction networks, etc. are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities. For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations; In a protein-protein interaction network, the nodes represent proteins and links represent their pairwise interactions. Two of the key problems presented by such data include node classification and link prediction (Li and Pi, 2020; Kumar et al., 2020). The state-of-the-art methods for solving such problems rely on graph representation learning methods (Zhou et al., 2022; Cai et al., 2018; Xu, 2021; Wang et al., 2022).
The state-of-the-art performance achieved by such methods comes at the expense of high computational costs – and the large carbon footprints and the attendant adverse environmental impacts (Desislavov et al., 2023; Strubell et al., 2020; de Vries, 2023; Thompson et al., 2021) – of data-hungry deep learning methods that rely on iterative parameter optimization and hyperparameter tuning. Against this background, we explore a computationally efficient, one-pass learning algorithm in which the model needs to see each data sample only once, as an alternative to state-of-the-art graph neural networks for node classification and link prediction.
Key Contributions: We introduce a Hyperdimensional Graph Learner (HDGL), a novel method for node classification and link prediction in graphs. HDGL maps node features into a very high-dimensional space (hyperdimensional or HD space for short) using the injectivity property of node representations in a family of Graph Neural Networks (GNNs) and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node yielding latent node representations that can support both node classification and link prediction tasks. HDGL provides a one-pass learning alternative to GNNs trained using computationally expensive iterative optimization methods. We report results of extensive experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL achieves accuracy that is competitive with that of the state-of-the-art GNN methods at substantially reduced computational cost; and on the link prediction task, HDGL matches the performance of DeepWalk and related methods, although it falls short of computationally demanding state-of-the-art GNN methods. We further show that HDGL, which does not require expensive iterative learning procedures, is well-suited for data/class-incremental learning, making it an attractive alternative to GNNs.
2. Related Work
@@
Graph Neural Networks (GNNs), introduced by Sperduti and Starita (1997); Frasconi et al. (1998); Gori et al. (2005); Scarselli et al. (2009), which learn embeddings of nodes in a graph by aggregating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning (Kipf and Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017), because of their superior performance across a broad spectrum of applications (Wu et al., 2020; Zhang et al., 2018; Zhou et al., 2022). However, as noted above, their state-of-the-art performance comes at a high computational cost and adverse environmental impact. In contrast, HDGL needs only a single pass through the training data, and hence, offers a computationally efficient, low energy footprint alternative to GNN for node classification and link prediction.
@@
Hyperdimensional computing (Kanerva, 1988, 2009; Neubert et al., 2019; Thomas et al., 2021) offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector (Kanerva, 1988, 2009). Simple operations, like element-wise additions and dot products (Kanerva, 1988; Neubert et al., 2019; Thomas et al., 2021), are used to perform computations on the encoded objects. Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware (Kleyko et al., 2023a, b). Hence, there is a growing interest in the use of HD computing in machine learning (Neubert et al., 2019; Kleyko et al., 2023b; Ge and Parhi, 2020; Morris et al., 2022).
@@
Hyperdimensional Graph Encodings. Poduval et al. (2022) and Nunes et al. (2022) have investigated hyperdimensional representations of graphs for similarity-based retrieval and graph classification. Nunes et al. (2022) constructs the Graph HD vector using node features derived using PageRank (Brin and Page, 1998). Poduval et al. (2022) assign random HD-vectors to nodes and considers them as node features to build a Graph HD vector. Thus, both methods ignore node features, whereas our focus is on methods that learn node-level representations for node classification and link prediction.
Hyperdimensional Node Embeddings. Recent work has explored hyperdimensional embeddings of nodes in a graph Kang et al. (2022); Li et al. (2023). Specifically, Kang et al. (2022) proposed RelHD, a novel processing-in-memory (PIM) hardware architecture (Mutlu et al., 2022) based on FeFET technology (Beyer et al., 2020). Li et al. (2023) introduced the HyperNode method for learning HD embeddings of nodes. However, both RelHD and HyperNode encode node features as binary indicator variables before mapping them to HD space and hence are unsuited for graphs with continuous or integer-valued node attributes e.g., word counts. Furthermore, none of the aforementioned methods (Nunes et al., 2022; Poduval et al., 2022; Kang et al., 2022; Li et al., 2023) support link prediction on graphs.
3. Preliminaries
We proceed to briefly summarize the key concepts needed to set the stage for introducing HDGL.
3.1. Node Classification and Link Prediction on Graphs
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of nodes, $\mathcal{E}$ represents the set of undirected edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$. Let $N=|\mathcal{V}|$ denote the number of nodes in the graph. Supposed each node $v\in\mathcal{V}$ is described by a tuple of $d$ features $\mathbf{x}_{v}\in\mathbb{R}^{d}$. The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$. The node classification problem can be described as follows: Given labels $y_{w}\in\{1,\cdots,L\}$ for each node $w\in\mathcal{W}$, where $\mathcal{W}\subset\mathcal{V}$; the task is to predict the labels of unlabeled nodes, i.e., nodes in $\mathcal{V}\setminus\mathcal{W}$. Similarly, the link prediction problem entails predicting the missing edges/new edges in $\mathcal{G}$, e.g., predicting new links/recommendation in a social network. We aim to solve both problems by learning a latent representation $\phi(v)$ for each node $v$, utilizing the structure of $\mathcal{G}$, i.e., the connectivity between nodes, the feature vector $\mathbf{x}_{v}$, and the node labels $y_{w}$.
3.2. Hyperdimensional Computing
@@
Hyper-dimensional computing, originally introduced by Kanerva (1988) is a brain-inspired approach to representing information that encodes each object $x\in\mathcal{X}$ using a high dimensional mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large (Kanerva, 1988, 2009). All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products (Kanerva, 1988; Neubert et al., 2019). The mapping $\phi$ is often random, and low precision which lends itself to fast, low-power, energy-efficient hardware realizations (Kleyko et al., 2023a, b). In this paper, we assume =$\mathcal{H}=\{0,1\}^{n}$. We proceed to summarize some of the key properties of HD representation.
@@
Near Orthogonality of Random vectors in HD Space. HD encodings of objects are produced by sampling each dimension independently and uniformly from a distribution. Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$. We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ (Kanerva, 1988). This is also known as a blessing of HD spaces.
@@
High Memory Capacity of HD Representation. As noted by Neubert et al. (2019), in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$. For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
@@
Noise or Error Tolerance. Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits (Ahmad and Hawkins, 2015; Neubert et al., 2019; Thomas et al., 2021). Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms (Neubert et al., 2019).
We next briefly describe some of the key operations used to compute with HD representations:-
@@
Bundling, denoted by $\oplus$, is a bitwise majority operation represented by :$\oplus:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. Bundling randomly sampled HD vectors results in a new vector which is similar to the input vectors. Bundling $\oplus$ is associative and commutative i.e $\mathbf{a}\oplus(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\oplus\mathbf{b})\oplus\mathbf{c}=(\mathbf{c}\oplus\mathbf{a})\oplus\mathbf{b}$. Note that in case of Bundling an even number of HD vectors, there is a possibility of ties since Bundling is a bitwise majority operation. To break ties, a tie-breaking policy needs to be used, e.g., random choice of 0 or 1 on dimensions where ties occur. Bundling can be thought of as the HD space analog of the mean operation euclidean spaces. Bundling, because it commutes, can be used to represent sets as well as multisets of objects encoded using HD vectors.
@@
Rotation, also called permutation, denoted by $\Pi$, is a bitwise left shift operation $\Pi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$. Rotation was introduced by Gayler (1998) in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$. In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
@@
3.3. Graph Neural Networks